My sister is having trouble with her math, and her teacher won't help her. I think this is a linear equation. In standard form already. It is confusing us.
The problem is
4x+y=0 Standard Form, y-intercept
A=4 B=1 C=0
Her teacher did
4x+1y=0
-4x -4x
y= -4/1+0 (The / here means over) for y-intercept she had m=-4/-1 b=0 HOw does this work? She doesn't believe in doing it any other way, but her way could someone please help us out here? Please explain, and another thing the teacher doesn't believe in subtraction.

Answer by jim_thompson5910(35256)

If you want to convert 4x+y = 0 to slope intercept form then

4x + y = 0

4x + y - 4x = 0 - 4x

y = -4x

The equation is now in slope intercept form y = mx+b where m = -4 (the slope) and b = 0 (the y-intercept)

--------------------------------------------------------------------------------------------------------------

If you need more help, you can email me at jim_thompson5910@hotmail.com

To find other ways to contact me, you can also visit my website: http://www.freewebs.com/jimthompson5910/home.html

Thanks,

Jim

--------------------------------------------------------------------------------------------------------------